The government spent $ 49 billion for various infrastructure upgrades. If the marginal propensity to consume is 0.74 ?, the marginal propensity to save is ? . ?(Enter your response rounded to two decimal? places.) The increased government spending created ?$ ? billion of additional spending. ?(Enter your response rounded to two decimal? places.)
Answer #1

MPC + MPS = 1

0.74 + MPS =1

MPS = 1- 0.74

= 0.26  

Govt multiplier  = 1/(1- MPC)

= 1/(1 - 0.74)

= 1/0.26

= 3.84  

so the increased government spending is = 3.84(49)  

= 188.46